# Euclid Quick Data Release (Q1)

### v1

@@
Strong gravitational lensing has the potential to provide a powerful probe of astrophysics and cosmology, but fewer than 1000 strong lenses have been confirmed previously. With resolution covering a third of the sky, the Euclid telescope will revolutionise strong lens finding, with lenses forecasted to be discovered amongst its 1.5 billion galaxies. We present an analysis of the performance of five machine-learning models at finding strong gravitational lenses in the quick release of Euclid data (Q1), covering 63 deg2. The models are validated with citizen scientists and expert visual inspection. We focus on the best performing network: a fine-tuned version of the Zoobot pretrained model, originally trained to classify galaxy morphologies in heterogeneous astronomical imaging surveys. Of the one million Q1 objects that Zoobot was tasked to find strong lenses within, the top 1000 ranked objects contained 122 grade A lenses (almost certain lenses), and 41 grade B lenses (probable lenses). A deeper search with the five networks combined with visual inspection discovered 250 (247) grade A (B) lenses, of which 224 (182) are ranked in the top by Zoobot. When extrapolated to the full Euclid survey, the highest ranked one million images will contain grade A or B strong gravitational lenses.
@@
Covering approximately deg2, the Euclid Wide Survey (EWS; Euclid Collaboration: Mellier et al. 2024) has the power to revolutionise strong gravitational lensing by galaxies, transforming it from a sample of about a thousand to over a hundred thousand (Collett 2015). Whilst many telescopes on the ground have surveyed large areas of sky, no previous wide area survey has ever had sufficient resolution to resolve multiple images of a single background source around a typical lens galaxy. The Einstein radius distribution of galaxy-galaxy lenses in the Universe is expected to peak at (Collett 2015), and many of these are within reach of Euclid with the full width at half maximum (FWHM) in the VIS filter (Euclid Collaboration: McCracken et al. 2025).
With exceptional resolution comes exceptional challenges of data volume. Euclid will observe around 1.5 billion unlensed galaxies (Euclid Collaboration: Scaramella et al. 2022) in its deg2, and manually searching these is intractable. Even the largest citizen science lens search projects have only inspected up to around one million images (Marshall et al. 2015; More et al. 2016; Sonnenfeld et al. 2020; González et al. 2025). To deliver the promise of Euclid strong lens finding, it is clear that automated methods must be employed to pre-screen candidates.
Recently, machine learning has gained substantial traction in the field of strong lens finding, with many teams developing machine-learning-based classifiers (Metcalf et al. 2019). Lanusse et al. (2018) and Avestruz et al. (2019) developed deep convolutional neural network (CNN) architectures that performed extremely well on simulated lenses (Metcalf et al. 2019), whilst Jacobs et al. (2017) trained a CNN to find lens candidates per square degree in the Canada–France–Hawaii Telescope (CFHT) data. CNNs have now been successfully employed to find lenses in the Dark Energy Survey (DES, Jacobs et al. 2019b, a; Rojas et al. 2022), the Kilo-Degree Survey (KiDS, Petrillo et al. 2017, 2019; Li et al. 2020, 2021; Nagam et al. 2023, 2024; Grespan et al. 2024), Panoramic Survey Telescope and Rapid Response System (PanSTARRS, Cañameras et al. 2020), the Ultraviolet Near-Infrared Optical Northern Survey (UNIONS, Savary et al. 2022), and Hyper Suprime-Cam (HSC) data (Shu et al. 2022; Jaelani et al. 2024). More recently González et al. (2025) showed that multi-class vision transformers are powerful alternatives to CNNs.
@@
The Euclid Quick Data Release 1 (Q1; Euclid Quick Release Q1 2025) data set gives the first opportunity to search for lenses in a large area (63 deg2) homogeneous data set with resolution. Prior to this, the only Euclid data available for finding lenses within has been the Euclid Early Release Observations (ERO, Euclid Early Release Observations 2024) data, which covered an area of 8.12 . With these data, an area of 0.7 was entirely visually inspected down to by experts where three grade A and 13 grade B lens candidates were found (Acevedo Barroso et al. 2024b). These lenses were used to test machine-learning models at finding lenses in Euclid (Pearce-Casey et al. 2024), and these models were then applied to the rest of the ERO, data where 14 grade A and 31 grade B strong lens candidates were found (Nagam et al. 2025).
This paper describes the efforts within the Euclid Consortium to train machine-learning models to discover strong lenses in Q1, which is part of the five paper series ‘The Strong Lensing Discovery Engine’. The series consists of: the overview paper Euclid Collaboration: Walmsley et al. (2025) that includes a description of all the lenses found in Q1; Euclid Collaboration: Rojas et al. (2025) that describes our effort to build a strong lens training set for Q1; this paper that describes the machine-learning lens finding effort; Euclid Collaboration: Li et al. (2025) that analyses the double-source-plane lens (DSPL) candidates that were discovered; and Euclid Collaboration: Holloway et al. (2025) that explores how to combine multiple lens-finding classifiers into an ensemble classifier for future lens searches.
The present paper is structured as follows. In Sect. 2 we discuss the data used in this work, which consists of the Euclid Q1 data (Sect. 2.1) and the machine-learning training data (Sect. 2.2). In Sect. 3 we describe the Zoobot machine-learning model and present what was learnt from optimising Zoobot to find lenses. In Sect. 4 we explore the other machine-learning models used in this work. The results are presented in Sect. 5, which include the results from a preliminary inspection of each model’s top 1000 ranked objects (Sect. 5.1), the results from combining machine learning with citizen science to find lenses in Q1 (Sect. 5.2), and a statistical analysis of how many lenses there are in Q1 beyond what was visually inspected (Sect. 5.3). A discussion of what was found is presented in Sect. 6 and the conclusions are presented in Sect. 7.
@@ 2 Data
Euclid collects data with two instruments, the visible imager (VIS; Euclid Collaboration: Cropper et al. 2024) and the near-infrared spectrometer and photometer (NISP; Euclid Collaboration: Jahnke et al. 2024). There are four bands in total, the VIS band and the NISP , , and bands. Although Euclid images are originally in FITS (Flexible Image Transport System) format, the quantity of data meant that it was necessary to work with JPEG (Joint Photographic Experts Group) files, which required a choice of scaling and filters. Three choices of filter combinations with two image scalings were chosen, producing a total of six combinations. The filter combinations used were -only greyscale images, + colour images, and + colour images. In the case of the colour images, the red/blue channels correspond to the longest/shortest wavelength images, and the green channel is the median of the two. All images are produced at the highest resolution of the Euclid bands, that of VIS , at per pixel (see Euclid Collaboration: Walmsley et al. 2025 for details). The two scalings chosen were , which provides a high contrast that amplifies both faint arcs and background noise, and the midtone-transfer function (MTF; see Euclid Collaboration: Walmsley et al. 2025), which produces a more homogenous image without saturating bright pixels. An example of these images for one cutout is shown in Fig. 1. Although converting images from FITS to JPEG files necessarily means some information is lost, the full 16 bits of information per pixel in FITS files results in large quantities of mostly redundant information. Providing that the scaling to the 8-bit format of JPEG images is such that it preserves most of the dynamic range, the amount of information lost is made up for by the significantly faster processing time. We find that for both the brightest and the faintest of lenses (whether these be the lenses we find in Q1 or lenses found in previous searches), the lenses are generally as easy to distinguish in the JPEG images as in the FITS images. These options were chosen based on the appearance of the Q1 data, but the same scalings were applied to simulations and other training and validation data to ensure homogeneity. All images have a size of pixels, corresponding to an angular scale of in the band. This size allows the full Einstein ring of all galaxy-scale lenses to be visible but does not necessarily encompass group-scale lenses.
@@ 2.1 Q1 data
Euclid collects data for two main surveys: the EWS and the Euclid Deep Survey (EDS). The former aims to survey around a third of the sky, covering the area that excludes the Galactic and ecliptic plane, amounting to , while the latter surveys three selected areas (Euclid Deep Field North, Euclid Deep Field South, and Euclid Deep Field Fornax), totalling , but with repeated exposures producing images that are up to 2 magnitudes deeper. The Euclid Q1 release covers the Deep Field areas, but at the depth of the EWS. Q1 provides the first look into what we will be able to find within Euclid’s lifetime and provides us with the first sample of Euclid lenses.
The Euclid data products we used were the background-subtracted merged mosaic tiles, in which the four images from the different Euclid bands are matched to the resolution of VIS . While has a PSF FWHM of (Euclid Collaboration: Cropper et al. 2024), for , , and the PSF FWHM is , , and , respectively (Euclid Collaboration: Jahnke et al. 2024), and the details of the interpolation are described in Euclid Collaboration: Romelli et al. (2025). Of all the objects in Q1, passed our selection criteria of , along with a quality check to remove objects such as stars (see Euclid Collaboration: Walmsley et al. 2025), and it was these objects that were analysed by the machine-learning models. Using cutouts with size means there will inevitably be overlap between images, but we ensure that the numbers of lenses reported has duplicate lenses removed.
@@ 2.2 Training and validation data sets
For the Q1 lens finding challenge, any member of the Euclid Collaboration strong lensing working group was invited to develop a machine-learning model. Many of the models used in this work were built on models developed for the strong lens search in the Euclid ERO data (Acevedo Barroso et al. 2024b; Pearce-Casey et al. 2024; Nagam et al. 2025), which were the first investigations of the lens-finding ability of Euclid. Given the aim of finding as many lenses as possible, there were few restrictions placed on how the models should be developed; rather, they were all developed independently and evaluated on their Q1 lens-finding performance.
@@
At the time of training the machine-learning models in preparation for Q1, there were multiple options of what images to use. The different data sets available consisted of simulated lenses, known real lenses, and non-lenses. The different machine-learning models were developed independently and as such were free to choose any available data sets for training and validation of the models. All of these data sets were available as JPEG files in all of the six combinations of scalings and filters, and it was left to the developers of each model to choose which of these to use for their final model. By allowing a diversity of options in terms of choice of training data and image processing, we are more likely to be able to find an optimal approach for lens finding in future Euclid data releases, allowing us to later refine these options to be able to make more direct comparisons between models. In the future when a large number of real Euclid lenses are available, we will be able to quantify different machine-learning approaches in a fully data-driven way.
@@ 2.2.1 Simulated lenses
The first set of simulations (S1) was based on the work presented in Euclid Collaboration: Rojas et al. (2025) and was created as follows. They simulated strong lensing systems using all four Euclid bands, following the procedure outlined in Rojas et al. (2022) and utilising the Lenstronomy package (Birrer & Amara 2018; Birrer et al. 2021). Deflectors were selected as LRGs with known redshifts and velocity dispersions, modelled using a Srsic profile in the band to estimate ellipticities and central positions. These parameters were optimised with a downhill simplex algorithm. Background sources were drawn from the Hubble Space Telescope Advanced Camera for Surveys (ACS) F814W high-resolution catalogue (Leauthaud et al. 2007; Scoville et al. 2007; Koekemoer et al. 2007), combined with HSC colour information (Cañameras et al. 2020). VIS magnitudes were computed by combining -band + -band images, while infrared magnitudes were matched to COSMOS 2020 VISTA (Visible and Infrared Survey Telescope for Astronomy) bands (Weaver et al. 2022). They paired lenses and sources, such that they produce an Einstein radius in the range –. A singular isothermal ellipsoid (SIE) mass model was adopted, with parameters derived from the Srsic fit and Einstein radii calculated using the redshifts and velocity dispersions of the lens and source. Source images were lensed based on the mass model, and then downsampled to match Euclid VIS image resolution. This last step includes convolution with the telescope point-spread function (PSF) and scaling to the lens image flux. Finally, the resulting simulated lensing features are added into the VIS and NISP lens images. Through augmenting the number of LRGs available for simulating by using four rotations of each LRG, the number of simulations was . There was an additional earlier version of the same simulations containing simulations, making a total number of simulations available.
@@
A second set of simulations (S2) was created using the GLAMER lensing code (Metcalf & Petkova 2014; Petkova et al. 2014). These simulations started by cutting out pixel images of all the observed galaxies with ¡ 22. Each image was matched to an object in the Flagship simulation (Euclid Collaboration: Castander et al. 2024) with a nearest neighbour algorithm in the space of all magnitudes in all four bands, ellipticity, and redshift, when available for the observed galaxy. The parameters of the Flagship galaxy and dark matter halo were then used to construct a mass model for the lens. The mass model consisted of a baryonic part in the form of a Srsic profile and a dark matter halo in the form of a Navarro–Frenk–White profile (Navarro et al. 1997). The position angle was set to that measured in the data. A source redshift was drawn from a distribution fit to the COSMOS 2020 FARMER catalogue -band number counts (Weaver et al. 2022). Rays were then shot through the mass profile to the source redshift. The caustics and critical curves were found. The lens was rejected if the Einstein radius was less than . The Einstein radius is defined as the radius of a circle that has the same area as the area within the tangential critical curve. A source was then selected and placed near or in the tangential caustic. The sources were modelled with multiple Srsic profiles. The image of the lensed object was then convolved with the local PSF, and Poisson noise was added. The simulations were performed at the resolution of VIS , then downsampled to the respective resolutions for VIS and NISP before being convolved with the PSFs. The NISP images were then interpolated to the higher VIS resolution using a BILINEAR interpolation kernel, similar to the processing of the real Euclid data. Another rejection step was taken after combining the lensed image with the original data image. Cases where the lensed image was too low in signal-to-noise or in contrast to the lens galaxy light were removed when visually inspecting the images. In these simulations, the mass of neighbouring objects was not included. See Metcalf et al. (2025. in prep.) for more details on these and related simulations. 5363 images were created without any augmentation.
@@
A third set of -only simulations (S3) was created, implementing the code presented by Schuldt et al. (2021). To create galaxy-scale lens simulations, a lens mass distribution modelled with an SIE profile was used, with the following input parameters: lens redshift, source redshift, and velocity dispersion. The lens centroid, axis ratio, and position angle were derived from the first and second brightness moments of the Euclid VIS images. Sources were randomly placed in the source plane, restricting the simulations to those with an Einstein radius between and . Subsequently, the sources were lensed in the image plane using the GLEE software (Suyu & Halkola 2010; Suyu et al. 2012). Finally, the simulated lensed source image was convolved with the Euclid PSF, rescaled according to Euclid s zero points, and combined with the original LRG image. The final simulations were given as images of , with simulations produced.111We note that due to the nature of the internal Euclid Collaboration Q1 timeline, this set of simulations was made available after many machine-learning models had started to be developed and hence was employed by fewer models.
@@
All of the simulations used for training employed painting simulated lensed sources on to real galaxies, as opposed to a fully simulated approach. We used three sets of simulations to increase training diversity. Two (S1 and S3) painted lensed sources on to galaxies with spectroscopically measured velocity dispersions, improving realism at the cost of using only bright LRGs as lenses (Euclid Collaboration: Rojas et al. 2025, Melo et al. in prep.). As high velocity dispersion LRGs were selected for this, the resulting simulations had disproportionately large Einstein radii, which could potentially bias the machine-learning models towards more massive lenses which are generally easier to detect. The other simulation set (S2) used proxies to avoid using spectra, resulting in potentially less physically realistic simulations but meaning that a much larger number of simulations could be produced around a wider range of galaxy types and masses (Metcalf 2025 in prep.). Only S1 and S2 produced simulations in the NISP bands. S1 produced the NISP images directly at the higher VIS resolution, while S2 produced the NISP simulations at the lower resolution and then interpolated this to the resolution of VIS , taking into account the correlated noise that exists in the real Euclid NISP images.

### v2

@@
Strong gravitational lensing has the potential to provide a powerful probe of astrophysics and cosmology, but fewer than 1000 strong lenses have been confirmed so far. With a resolution covering a third of the sky, the Euclid telescope will revolutionise the identification of strong lenses, with lenses forecasted to be discovered amongst the 1.5 billion galaxies it will observe. We present an analysis of the performance of five machine-learning models at finding strong gravitational lenses in the quick release of Euclid data (Q1) covering 63 deg2. The models have been validated by citizen scientists and expert visual inspection. We focus on the best-performing network: a fine-tuned version of the Zoobot pretrained model originally trained to classify galaxy morphologies in heterogeneous astronomical imaging surveys. Of the one million Q1 objects that Zoobot was tasked to find strong lenses within, the top 1000 ranked objects contain 122 grade A lenses (almost-certain lenses) and 41 grade B lenses (probable lenses). A deeper search with the five networks combined with visual inspection yielded 250 (247) grade A (B) lenses, of which 224 (182) are ranked in the top by Zoobot. When extrapolated to the full Euclid survey, the highest ranked one million images will contain grade A or B strong gravitational lenses.
@@
Covering approximately deg2, the Euclid Wide Survey (EWS; Euclid Collaboration: Mellier et al. 2025) has the power to revolutionise strong gravitational lensing by galaxies: the strong lenses Euclid is forecasted to discover will provide a sample two orders of magnitude larger than what was previously known (Collett 2015). Whilst many telescopes on the ground have surveyed large areas of sky, no previous wide area survey has ever had sufficient resolution to resolve multiple images of a single background source around a typical lens galaxy. The Einstein radius distribution of galaxy-galaxy lenses in the Universe is expected to peak at (Collett 2015), and many of these lenses are within reach of Euclid with its full width at half maximum (FWHM) in the filter (Euclid Collaboration: McCracken et al. 2025).
With exceptional resolution comes exceptional challenges of data volume. Euclid will observe around 1.5 billion unlensed galaxies (Euclid Collaboration: Scaramella et al. 2022) in its deg2, and manually searching these is intractable. Even the largest citizen science lens search projects have only inspected up to around one million images (Marshall et al. 2015; More et al. 2016; Sonnenfeld et al. 2020; González et al. 2025). To deliver on the promise of Euclid’s strong lens finding capabilities, it is clear that automated methods must be employed to pre-screen candidates.
Recently, machine learning has gained substantial traction in the field of strong lens finding, with many teams developing machine-learning-based classifiers (Metcalf et al. 2019). Lanusse et al. (2018) and Avestruz et al. (2019) developed deep convolutional neural network (CNN) architectures that perform extremely well on simulated lenses (Metcalf et al. 2019), whilst Jacobs et al. (2017) trained a CNN to find lens candidates per square degree in the Canada–France–Hawaii Telescope (CFHT) data. CNNs have now been successfully employed to find lenses in the Dark Energy Survey (DES; Jacobs et al. 2019b, a; Rojas et al. 2022), the Dark Energy Spectroscopic Instrument (DESI) Legacy Survey (Huang et al. 2020, 2021; Storfer et al. 2024), the Kilo-Degree Survey (KiDS; Petrillo et al. 2017, 2019; Li et al. 2020, 2021; Nagam et al. 2023, 2024; Grespan et al. 2024), Panoramic Survey Telescope and Rapid Response System (PanSTARRS; Cañameras et al. 2020), the Ultraviolet Near-Infrared Optical Northern Survey (UNIONS; Savary et al. 2022), Hyper Suprime-Cam (HSC; Shu et al. 2022; Jaelani et al. 2024), and Hubble Space Telescope (HST) data (Pourrahmani et al. 2018). More recently, González et al. (2025) showed that multi-class vision transformers are powerful alternatives to CNNs.
@@
The Euclid Quick Data Release 1 (Q1; Euclid Quick Release Q1 2025) data set provides the first opportunity to search for lenses in a homogeneous data set covering a large area (63 deg2) with resolution. Prior to this, the only Euclid data available for finding lenses has been the Euclid Early Release Observations (ERO; Euclid Early Release Observations 2024) data, which covered an area of 8.12 . With these data, an area of 0.7 was entirely visually inspected down to by experts, and three grade A and 13 grade B lens candidates were found (Acevedo Barroso et al. 2025b). These lenses were used to test machine-learning models at finding lenses in Euclid (Pearce-Casey et al. 2025), and these models were then applied to the rest of the ERO data, resulting in 14 grade A and 31 grade B strong lens candidates being found (Nagam et al. 2025).
This paper describes the efforts within the Euclid Consortium to train machine-learning models to discover strong lenses in Q1, which is part of the five-paper series ‘The Strong Lensing Discovery Engine’. The series consists of the overview paper Euclid Collaboration: Walmsley et al. (2025a), which includes a description of all the lenses found in Q1; Euclid Collaboration: Rojas et al. (2025), which describes our effort to build a strong lens training set for Q1; this paper, which describes the machine-learning lens finding effort; Euclid Collaboration: Li et al. (2025), which analyses the double-source-plane lens (DSPL) candidates that were discovered; and Euclid Collaboration: Holloway et al. (2025), which explores how to combine multiple lens-finding classifiers into an ensemble classifier for future lens searches.
The present paper is structured as follows. In Sect. 2 we discuss the data used in this work, which consists of the Euclid Q1 data (Sect. 2.1) and the machine-learning training data (Sect. 2.2). In Sect. 3 we describe the Zoobot machine-learning model and present what was learnt from optimising Zoobot to find lenses. In Sect. 4 we explore the other machine-learning models used in this work. The results are presented in Sect. 5, which include the results from a preliminary inspection of each model’s top 1000 ranked objects (Sect. 5.1), the results from combining machine learning with citizen science to find lenses in Q1 (Sect. 5.2), and a statistical analysis of how many lenses there are in Q1 beyond the sample that was visually inspected (Sect. 5.3). A discussion of what was found is presented in Sect. 6, and the conclusions are presented in Sect. 7.
@@ 2 Data
Euclid collects data with two instruments, the visible imager (VIS; Euclid Collaboration: Cropper et al. 2025) and the near-infrared spectrometer and photometer (NISP; Euclid Collaboration: Jahnke et al. 2025). There are four bands in total, the VIS band and the NISP , , and bands. Although Euclid images are originally in Flexible Image Transport System (FITS) format, the quantity of data meant that it was necessary to work with JPEG (Joint Photographic Experts Group) files, which required a choice of scaling and filters. Three choices of filter combinations with two image scalings were chosen, producing a total of six combinations. The filter combinations used were -only greyscale images, + colour images, and + colour images. In the case of the colour images, the red (blue) channels correspond to the longest (shortest) wavelength images, and the green channel is the median of the two. All images have been produced at the highest resolution of the Euclid bands, that of VIS , at per pixel (see Euclid Collaboration: Walmsley et al. 2025a for details). The two scalings chosen were , which provides a high contrast that amplifies both faint arcs and background noise, and the midtone-transfer function (MTF; see Euclid Collaboration: Walmsley et al. 2025a), which produces a more homogeneous image without saturating bright pixels. An example of these images for one cutout is shown in Fig. 1. Although converting images from FITS to JPEG files necessarily means some information is lost, the full 16 bits of information per pixel in FITS files results in large quantities of mostly redundant information. Providing that the scaling to the 8-bit format of JPEG images is such that it preserves most of the dynamic range, the amount of information lost is made up for by the significantly faster processing time. We find that for both the brightest and the faintest of lenses (whether these be the lenses we find in Q1 or lenses found in previous searches), the lenses are generally as easy to distinguish in the JPEG images as in the FITS images. These options were chosen based on the appearance of the Q1 data, but the same scalings were applied to simulations and other training and validation data to ensure homogeneity. All images have a size of pixels, corresponding to an angular scale of in the band. This size allows the full Einstein ring of all galaxy-scale lenses to be visible but does not necessarily encompass group-scale lenses.
@@ 2.1 Q1 data
Euclid collects data for two main surveys: the EWS and the Euclid Deep Survey (EDS). The former aims to survey around a third of the sky, covering the area that excludes the Galactic and ecliptic plane, amounting to , while the latter surveys three selected areas (Euclid Deep Field North, Euclid Deep Field South, and Euclid Deep Field Fornax) totalling , but with repeated exposures producing images that are up to two magnitudes deeper. The Euclid Q1 release covers the Deep Field areas but at the depth of the EWS. Q1 provides the first look into what we will be able to find within Euclid’s lifetime and provides us with the first sample of Euclid lenses.
The Euclid data products we used were the background-subtracted merged mosaic tiles, in which the four images from the different Euclid bands are matched to the resolution of VIS . While has a PSF FWHM of (Euclid Collaboration: Cropper et al. 2025), for , , and the PSF FWHM is , , and , respectively (Euclid Collaboration: Jahnke et al. 2025) and the details of the interpolation are described in Euclid Collaboration: Romelli et al. (2025). Of all the objects in Q1, passed our selection criteria of , which was motivated by the findings from the lens search in the ERO data (Acevedo Barroso et al. 2025b). There was an additional quality check to remove objects such as stars (for further details see Euclid Collaboration: Walmsley et al. 2025a). It was these remaining objects that were analysed by the machine-learning models. Using cutouts with size means there will inevitably be overlap between images, but we ensure that the numbers of lenses reported has duplicate lenses removed.
@@ 2.2 Training and validation data sets
For the Q1 lens finding challenge, any member of the Euclid Collaboration strong lensing working group was invited to develop a machine-learning model. Many of the models used in this work were built on models developed for the strong lens search in the Euclid ERO data (Acevedo Barroso et al. 2025b; Pearce-Casey et al. 2025; Nagam et al. 2025), which were the first investigations of the lens-finding ability of Euclid. Given the aim of finding as many lenses as possible, there were few restrictions placed on how the models should be developed; rather, they were all developed independently and evaluated on their Q1 lens-finding performance.
@@
At the time of training the machine-learning models in preparation for Q1, there were multiple options of what images to use. The different data sets available consisted of simulated lenses, known real lenses, and non-lenses. The different machine-learning models were developed independently and as such were free to choose any available data sets for training and validation of the models. All of these data sets were available as JPEG files in all the six combinations of scalings and filters, and it was left to the developers of each model to choose which of these to use for their final model. By allowing a diversity of options in terms of choice of training data and image processing, we are more likely to be able to find an optimal approach for lens finding in future Euclid data releases, allowing us to later refine these options to be able to make more direct comparisons between models. In the future when a large number of real Euclid lenses are available, we will be able to quantify different machine-learning approaches in a fully data-driven way.
@@ 2.2.1 Simulated lenses
The first set of simulations (S1) was based on the work presented in Euclid Collaboration: Rojas et al. (2025) and was created as follows. They simulated strong lensing systems using all four Euclid bands, following the procedure outlined in Rojas et al. (2022) and utilising the Lenstronomy package (Birrer & Amara 2018; Birrer et al. 2021). Deflectors were selected as LRGs with known redshifts and velocity dispersions, modelled using a Srsic profile in the band to estimate ellipticities and central positions. These parameters were optimised with a downhill simplex algorithm. Background sources were drawn from the HST Advanced Camera for Surveys F814W high-resolution catalogue (Leauthaud et al. 2007; Scoville et al. 2007; Koekemoer et al. 2007), combined with HSC colour information (Cañameras et al. 2020). VIS magnitudes were computed by combining -band + -band images, while infrared magnitudes were matched to COSMOS 2020 VISTA (Visible and Infrared Survey Telescope for Astronomy) bands (Weaver et al. 2022). They paired lenses and sources such that they produce an Einstein radius in the range –. A singular isothermal ellipsoid (SIE) mass model was adopted, with parameters derived from the Srsic fit and Einstein radii calculated using the redshifts and velocity dispersions of the lens and source. Source images were lensed based on the mass model, and then downsampled to match Euclid VIS image resolution. This last step included convolution with the telescope point-spread function (PSF) and scaling to the lens image flux. Finally, the resulting simulated lensing features were added into the VIS and NISP lens images. The simulation pipeline did not include a deconvolution with HST PSF. This may produce a small impact in the smoothness and shape of the simulated arc that under Euclid PSF would be negligible, and would never result in unrealistic lensing configurations. Through augmenting the number of LRGs available for simulating by using four rotations of each LRG, the number of simulations was . There was an additional earlier version of the same simulations containing simulations, making a total number of simulations available.
@@
A second set of simulations (S2) was created using the GLAMER lensing code (Metcalf & Petkova 2014; Petkova et al. 2014). These simulations started by cutting out pixel images of all the observed galaxies with ¡ 22, with some additional cuts to remove stars and reduce the number of face on spirals. The resulting sample of galaxies exhibited a range of morphologies, including elliptical and disc galaxies, and reflect the diversity of galaxy morphologies in the Universe. Each image was matched to an object in the Flagship simulation (Euclid Collaboration: Castander et al. 2025) with a nearest-neighbour algorithm in the space of all magnitudes in all four bands, ellipticity, and redshift, when available for the observed galaxy. The parameters of the Flagship galaxy and dark matter halo were then used to construct a mass model for the lens. The mass model consisted of a baryonic part in the form of a Srsic profile and a dark matter halo in the form of a Navarro–Frenk–White profile (Navarro et al. 1997). The position angle was set to that measured in the data. A source redshift was drawn from a distribution fit to the COSMOS 2020 FARMER catalogue -band number counts (Weaver et al. 2022). Rays were then shot through the mass profile to the source redshift. The caustics and critical curves were found. The lens was rejected if the Einstein radius was less than . A source was then selected and placed near or in the tangential caustic. The source’s surface brightness was represented by between one and four Srsic profiles. To emulate some of the complex morphologies of high-redshift sources, the number of Srsic profiles, their relative brightnesses, and their positions were generated randomly from distributions tuned by eye. Their total brightnesses and overall sizes were based on randomly selected sources from the Hubble Ultra Deep Field with similar redshifts (see Meneghetti et al. 2008, 2010 for a discussion of this sample). The source redshifts were set to five, allowing the distribution of lens redshifts and mass parameters as well as source sizes and brightnesses to provide the variation in lens properties. The image of the lensed object was then convolved with the local PSF, and Poisson noise was added. The simulations were performed at the resolution of VIS , then downsampled to the respective resolutions for VIS and NISP before being convolved with the PSFs. The NISP images were then interpolated to the higher VIS resolution using a BILINEAR interpolation kernel, similar to the processing of the real Euclid data. Another rejection step was taken after combining the lensed image with the original data image. Cases where the lensed image was too low in signal-to-noise or in contrast to the lens galaxy light were removed when visually inspecting the images. In these simulations, the mass of neighbouring objects was not included. See Metcalf et al. (2025. in prep.) for more details on these and related simulations. 5363 images were created without any augmentation.
@@
A third set of -only simulations (S3) was created, implementing the code presented by Schuldt et al. (2021). To create galaxy-scale lens simulations, the same LRGs as in simulation set S1 were used. The lens mass distribution was modelled with an SIE profile, with the following input parameters: lens redshift, source redshift, and velocity dispersion. The lens centroid, axis ratio, and position angle were derived from the first and second brightness moments of the Euclid VIS images. The source galaxies originated from the catalogue presented in Schuldt et al. (2021), using images of high-redshift galaxies from the Hubble Ultra Deep Field (Inami et al. 2017). Sources were randomly placed in the source plane, restricting the simulations to those with an Einstein radius between and . Subsequently, the sources were lensed in the image plane using the GLEE software (Suyu & Halkola 2010; Suyu et al. 2012). Finally, the simulated lensed source image was convolved with the Euclid PSF, rescaled according to Euclid s zero points, and combined with the original LRG image. The final simulations were given as images of , with simulations produced.111We note that due to the nature of the internal Euclid Collaboration Q1 timeline, this set of simulations was made available after many machine-learning models had started to be developed and hence was employed by fewer models.
@@
All of the simulations used for training employed an approach of painting simulated lensed sources on to real galaxies, as opposed to a fully simulated approach. We used three sets of simulations to increase training diversity. The distributions of redshifts of the lenses and sources from the simulations are shown in Fig. 2 and Fig. 3 respectively, where they are compared to the distributions predicted by Collett (2015). These distributions mostly align for the lens redshifts, while there is much more variation in source redshift distribution. Two simulation sets (S1 and S3) painted lensed sources on to galaxies with spectroscopically measured velocity dispersions, improving realism at the cost of using only bright LRGs as lenses (Euclid Collaboration: Rojas et al. 2025, Melo et al. in prep.). As high velocity dispersion LRGs were selected for this, the resulting simulations had disproportionately large Einstein radii. The distributions of Einstein radii of the simulations are compared to the results from the automated modelling of the Q1 lenses and the forecast by Collett (2015) in Fig. 4. Simulations S1 and S3 tend to produce larger Einstein radii than S2, and all simulation sets have larger radii than predicted by the forecast. The distribution from Q1 lens modelling aligns reasonably well with the simulations but is shifted to larger radii compared to the forecast. This suggests that the simulations could be biasing the machine learning models towards detecting lenses with larger Einstein radii. However, there are the additional effects that humans are better at recognising lenses with larger Einstein radii (Euclid Collaboration: Walmsley et al. 2025b), and that automated lens modelling may be more likely to succeed for such systems, both of which could also contribute to a bias in the distribution.
The simulation set S2 used proxies to avoid using spectra, resulting in potentially less physically realistic simulations but meaning that simulations could be produced around a wider range of galaxy morphologies and masses without being constrained by spectroscopic data availability (Metcalf 2025 in prep.). Only S1 and S2 produced simulations in the NISP bands. S1 produced the NISP images directly at the higher VIS resolution, while S2 produced the NISP simulations at the lower resolution and then interpolated this to the resolution of VIS , taking into account the correlated noise that exists in the real Euclid NISP images.